# On the verification of a Nicolas inequality

### v6

@@ 1. Introduction and notation
Nicolas inequality was introduced in [7, Theorem 2] and establish that one and only one of the following holds:
@@ 2. On the sign of the error term
In this unique section we prove a theorem and discuss the error term. The first one relates the sign of the error term of Mertens theorem for the sum of reciprocal of primes with the verification of the Nicolas inequality. To this aim we apply results involving the zeta prime function, basic estimates for Chevyshev functions and standard techniques of real analysis. In the second part we analyze the sign of such an error term, to this end we revise the proof of Mertens Theorem and obtain an explicit formula for the term, then we approximate sufficiently its value.
@@
Taking into account Mertens Theorem (see e.g. [8, Theorem 8.8(d)]) and the above equation, we can rewrite the Nicolas inequality, namely $e^{\gamma}\log\log N_{x}<\dfrac{N_{x}}{\varphi_{x}}$, as
@@
where $\gamma$ is the Euler constant and $b$ is known as the Mertens constant ([10, (2.7)]).
@@
Therefore $x^{e^{C/\log x}}>x(1+\frac{\alpha}{\log x})$ and for a suitable $\alpha$ we can ensure that $x^{e^{C/\log x}}>\theta(x)$ for every $x\geq x_{2}$, we can choose for instance $\alpha=1/2$ (see [10, Theorem 4(3.15)]). In consequence, taking logarithms twice at both sides, we first get $e^{C/\log x}\log x>\log\theta(x)$, and then $\log\Big(\frac{\log x}{\log\theta(x)}\Big)+\frac{C}{\log x}>0$ for every $x\geq x_{2}$. This implies that there exists some $\varepsilon>0$ such that
@@
Therefore Nicolas inequality holds for every $x\geq x_{3}$, which is not exactly (T), but it is in contradiction with (F), and as by [7, Theorem 2] there are only two possibilities, (T) must hold.
@@
Observe that the failure of Nicolas inequality depends directly from the change of sign of $\theta(x)-x$. In the following subsection we have used the quick guide of Mertens Theorem given in [2].
@@ 2.1. Discussion on the sign of the error term
Step 1. Let start with an estimate for the sum of logarithms (see [8, Lemma 8.2 and Theorem 8.8(a)] and also [2]).
$\sum_{n\leq x}\log x=x\log x-x+O_{1}(\log x)\,$,
@@
Step 2. We follow with a sum involving the Mangoldt function (see [8, Theorem 8.8(a)]).
@@
for $k\in(10^{4},10^{5})$ and a list of primes $\{p_{j}\}_{j}$. Alternatively we found estimates from above and from below for the second term. We can assume that $x$ is integer, hence the reminder of $\dfrac{x}{d}$ is $1$ for $d\in(x/2,x)$, so $\Big\{\dfrac{x}{d}\Big\}=\dfrac{1}{d}$ for $d\in(x/2,x)$, also recall that $\Big\{\dfrac{x}{x}\Big\}=0$. Analogously for $d\in(x/3,x/2)$, the reminder is $2$, so in this case $\Big\{\dfrac{x}{d}\Big\}=\dfrac{2}{d}$ and again $\Big\{\dfrac{x/2}{x/2}\Big\}=0$. Taking into account [8, Theorem 8.3], we can assert that $\psi(x/2)<3/4\psi(x)$ and $\psi(x/3)<1/2\psi(x)$ for a sufficient large integer $x$, thus
@@
In consequence, taking into account that $ax<\psi(x)<cx$ for $a=0.780355$... and $c=3/2a$, for a sufficiently large $x$ (see [8, Theorem 8.3]), we deduce that
@@
Step 3. Next we deal with a sum over primes (see [8, Theorem 8.8(b)]).
@@
Let us obtain two coarse bounds. We know that the second term is bounded and, since $\log p/p^{n}$ is positive, it increases when the $x$ grows and the partial sums in $n$ grows. Moreover it is bounded from above by $\sum_{j=2}^{\infty}\frac{\log j}{j(j-1)}$ (see proof of [8, Theorem 8.8(b)]), which is bounded by the integral of $\dfrac{\log(x-1)}{(x-1)(x-2)}$ between $2$ and infinity. So, for sufficiently large $x$
@@
Step 4. Finally we study the error term of the sum of reciprocal of primes (see [8, Theorem 8.8(d)]). We define $R(x)=\displaystyle\sum_{p\leq x}\dfrac{\log p}{p}-\log x$, so thanks to the estimate given in previous steps we get $R(x)=O_{3}(1)$. Let
@@
Estimates above are trivial, but we have seen that $R(x)$ is decreasing almost everywhere (a.e.) and negative for sufficiently large $x$. Let the expression of $b$ given in proof of [3, Theorem 427] (see also [2, Propostion]), that is
@@ Conclusion
Observe that $O(1/\log x)$ converges to zero and, from the last equation, it is the difference between an step function with discontinuities on primes and an strictly monotone function. We already saw this kind of functions in [9] (for instance), namely $\pi(x)-\operatorname{li}(x)$, $\psi(x)-x$ and $\theta(x)-x$, and we know they change of sign infinitely asymptoticaly. So it seems reasonable to think that $O(1/\log x)$ also change of sign intinitely for sufficiently large $x$, but this is another issue we do not deal here (see [4, Chapter V] for more information on these irregularities).
@@ Acknowledgment
We would like to thank Professors Wladyslaw Narkiewicz and Enrique A. Sánchez Pérez for their comments and suggestions.

### v7

@@ 1. Introduction and notation
Nicolas inequality was introduced in [8, Theorem 2] and establish that one and only one of the following holds:
@@ 2. On the sign of the error term
In this unique section we prove a theorem and discuss the error term. The first one relates the sign of the error term of Mertens theorem for the sum of reciprocal of primes with the verification of the Nicolas inequality. To this aim we apply results involving the zeta prime function, basic estimates for Chevyshev functions and standard techniques of real analysis. In the second part we analyze the sign of such an error term, to this end we revise the proof of Mertens Theorem and obtain an explicit formula for the term, then we approximate sufficiently its value. Unfortunately, the conditions we impose never hold by virtue of [1, Theorem 1.1], since the error term change of sign infinitely often, see also [11]. After that, we reformulate the theorem.
@@
Taking into account Mertens Theorem (see e.g. [9, Theorem 8.8(d)]) and the above equation, we can rewrite the Nicolas inequality, namely $e^{\gamma}\log\log N_{x}<\dfrac{N_{x}}{\varphi_{x}}$, as
@@
where $\gamma$ is the Euler constant and $b$ is known as the Mertens constant ([12, (2.7)]).
@@
Therefore $x^{e^{C/\log x}}>x(1+\frac{\alpha}{\log x})$ and for a suitable $\alpha$ we can ensure that $x^{e^{C/\log x}}>\theta(x)$ for every $x\geq x_{2}$, we can choose for instance $\alpha=1/2$ (see [12, Theorem 4(3.15)]). In consequence, taking logarithms twice at both sides, we first get $e^{C/\log x}\log x>\log\theta(x)$, and then $\log\Big(\frac{\log x}{\log\theta(x)}\Big)+\frac{C}{\log x}>0$ for every $x\geq x_{2}$. This implies that there exists some $\varepsilon>0$ such that
@@
Therefore Nicolas inequality holds for every $x\geq x_{3}$, which is not exactly (T), but it is in contradiction with (F), and as by [8, Theorem 2] there are only two possibilities, (T) must hold.
@@
Remark 2.2.
Notice that conditions of theorem above are too restrictive, in the sense that condition of the first part never holds taking into account equation [12, (2.3)], that is
$\sum_{p\leq x}\frac{1}{p}=\log\log x+b+O(\exp(-\alpha\sqrt{\log x}))\,$,
for some positive constant $\alpha$. In fact
Observe that the failure of Nicolas inequality depends directly from the change of sign of $\theta(x)-x$. From a revision of proof above we can deduce a
Corollary 2.3.
Let $x\geq 2$, and let the Mertens formula
(5) $\sum_{p\leq x}\frac{1}{p}=\log\log x+b+O(1/\log x)\,.$
(I) Assume that $\theta(x)>x$. If $O(1/\log x)<0$, then (1) does not hold for $x$.
(II) Assume that $\theta(x)\leq x$. If $O(1/\log x)\geq 0$, then (1) holds for $x$.
An alternative to theorem above is
Theorem 2.4.
Let $x>2$, and let the Mertens formula
(6) $\sum_{p\leq x}\frac{1}{p}=\log\log x+b+O(1/\log x)\,.$
We define
$B(x):=\sum_{p>x}\log\big(1-\dfrac{1}{p}\big)+\dfrac{1}{p}\,$,
(I) Assume that $\theta(x)>x$. If (1) holds for $x$, then $O(1/\log x)\geq-B(x)$.
(II) Assume that $\theta(x)\leq x$. If (1) does not hold for $x$, then $O(1/\log x)\leq-B(x)\,.$
Proof.
We deduce (I) and (II) from a the following reasonings. By rewriting proof of [9, Theorem 8.8(d)] we see that
$\prod_{p\leq x}\big(1-\frac{1}{p}\big)=\frac{e^{-\gamma}}{\log x}\exp(-(O(1/\log x)+B(x)))\,$,
where $O(1/\log x)$ is in (6), is the error term of Mertens Theorem and
$B(x)=\sum_{p>x}\log\big(1-\dfrac{1}{p}\big)+\dfrac{1}{p}\,$,
it is an exercise to show that $B(x)<0$. In our notation this can be written as
$C_{x}:=\frac{N_{x}}{\varphi_{x}\log x}=e^{\gamma}\exp(O(1/\log x)+B(x))=e^{\gamma}+\varepsilon_{x}\,$,
for some $\varepsilon_{x}\in(-1,1)$ for every $x>2$. We assumed that $x>2$ is fixed.
(I) We first assume that $\theta_{x}=\log N_{x}>x$, then taking logarithms at both sides and multiplying by $C_{k}$ we obtain
$C_{x}\log\log N_{x}>\frac{N_{x}}{\varphi_{x}}\,,\quad\mbox{and hence }\quad(e^{\gamma}+\varepsilon_{x})\log\log N_{x}>\frac{N_{x}}{\varphi_{x}}\,.$
In consequence, by assuming that (1) holds for x
$\varepsilon_{x}\log\log N_{x}>\frac{N_{x}}{\varphi_{x}}-e^{\gamma}\log\log N_{x}>0\,$,
since (1) holds. So $\varepsilon_{x}>0$, thus $C_{x}>e^{\gamma}$. From defintion of $C_{x}$ above we deduce that $\exp(O(1/\log x)+B(x))>1$, and so $O(1/\log x)>-B(x)>0$. In fact, if we assume that $O(1/\log x)\leq 0$, we get
$\log\Big(\frac{\log x}{\log\theta(x)}\Big)+O(1/\log x)\leq 0\,$,
and following the arguments of Step 3 in the proof of previous theorem we obtain
$\gamma\geq\gamma+\log\Big(\frac{\log x}{\log\theta(x)}\Big)+O(1/\log x)=\log\Big(\frac{N_{x}}{\varphi_{x}\log\log N_{x}}\Big)\,$,
hence Nicolas inequality does not hold for $x$.
(II) Now we assume that $\theta_{x}\leq x$. By the same argumentation as above, if we assume that (1) does not hold for $x$, we obtain in this case that $O(1/\log x)\leq-B(x)$, where B(x)-B(x) is positive.
Remark 2.5.
Observe that by definition of $B(x)$ and by Mertens Theorem the relation between $O(1/\log x)\lesseqgtr-B(x)$ is equivalent to
$\sum_{p\leq x}\frac{1}{p}\lesseqgtr b+\log\log x-\sum_{p>x}\log\big(1-\dfrac{1}{p}\big)+\dfrac{1}{p}=b+\log\log x+\sum_{p>x}\log\big(\dfrac{p}{p-1}\big)-\sum_{p>x}\dfrac{1}{p}\,.$
These inequalities could bring us to other discussions.
In the following subsection we have used the quick guide of Mertens Theorem given in [3].
@@ 2.1. Discussion on the sign of the error term
Step 1. Let start with an estimate for the sum of logarithms (see [9, Lemma 8.2 and Theorem 8.8(a)] and also [3]).
$\sum_{n\leq x}\log n=x\log x-x+O_{1}(\log x)\,$,
@@
Step 2. We follow with a sum involving the Mangoldt function (see [9, Theorem 8.8(a)]).
@@
for $k\in(10^{4},10^{5})$ and a list of primes $\{p_{j}\}_{j}$. Alternatively we found estimates from above and from below for the second term. We can assume that $x$ is integer, hence the reminder of $\dfrac{x}{d}$ is $1$ for $d\in(x/2,x)$, so $\Big\{\dfrac{x}{d}\Big\}=\dfrac{1}{d}$ for $d\in(x/2,x)$, also recall that $\Big\{\dfrac{x}{x}\Big\}=0$. Analogously for $d\in(x/3,x/2)$, the reminder is $2$, so in this case $\Big\{\dfrac{x}{d}\Big\}=\dfrac{2}{d}$ and again $\Big\{\dfrac{x/2}{x/2}\Big\}=0$. Taking into account [9, Theorem 8.3], we can assert that $\psi(x/2)<3/4\psi(x)$ and $\psi(x/3)<1/2\psi(x)$ for a sufficient large integer $x$, thus
@@
In consequence, taking into account that $ax<\psi(x)<cx$ for $a=0.780355$... and $c=3/2a$, for a sufficiently large $x$ (see [9, Theorem 8.3]), we deduce that
@@
Step 3. Next we deal with a sum over primes (see [9, Theorem 8.8(b)]).
@@
Let us obtain two coarse bounds. We know that the second term is bounded and, since $\log p/p^{n}$ is positive, it increases when the $x$ grows and the partial sums in $n$ grows. Moreover it is bounded from above by $\sum_{j=2}^{\infty}\frac{\log j}{j(j-1)}$ (see proof of [9, Theorem 8.8(b)]), which is bounded by the integral of $\dfrac{\log(x-1)}{(x-1)(x-2)}$ between $2$ and infinity. So, for sufficiently large $x$
@@
Step 4. Finally we study the error term of the sum of reciprocal of primes (see [9, Theorem 8.8(d)]). We define $R(x)=\displaystyle\sum_{p\leq x}\dfrac{\log p}{p}-\log x$, so thanks to the estimate given in previous steps we get $R(x)=O_{3}(1)$. Let
@@
Estimates above are trivial, but we have seen that $R(x)$ is decreasing almost everywhere (a.e.) and negative for sufficiently large $x$. Let the expression of $b$ given in proof of [4, Theorem 427] (see also [3, Propostion]), that is
@@ Conclusion
Observe that $O(1/\log x)$ converges to zero and, from the last equation, it is the difference between an step function with discontinuities on primes and an strictly monotone function. We already saw this kind of functions in [10] (for instance), namely $\pi(x)-\operatorname{li}(x)$, $\psi(x)-x$ and $\theta(x)-x$, and we know they change of sign infinitely asymptoticaly. So it seems reasonable to think that $O(1/\log x)$ also change of sign intinitely for sufficiently large $x$, but this is another issue we do not deal here (see [5, Chapter V] for more information on these irregularities).
@@ Acknowledgment
We would like to thank Professors Wladyslaw Narkiewicz, particularly for Remark 2.2, and Enrique A. Sánchez Pérez for their comments and suggestions.